OpenStudy (anonymous):

If Michael gave 1/12 of his money to Sally and 1/5 of his remaining money to Heidi. If he still had $704, how much money did he have originally?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

let M be the money he had originally....

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so... [M - (M/2)] - (1/5)[M - (M/2)] = 704

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

can't I backsolve the problem?

OpenStudy (anonymous):

\(\Large [M - (\color {red}{M/2})] - \color {blue}{(1/5)[M - (M/2)]} = 704 \) the red quantity is what he gave to sally... the blue quantity is what he gave to Heidi..... after giving away that money, he still has 704 left....

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

so

OpenStudy (anonymous):

yeah... i guess u could back track to solve this but i think doing the algebra should accomplish the same task... solve for M...

OpenStudy (anonymous):

I got M-M/2-1/5[M-m/2]=704

OpenStudy (anonymous):

M/2-M/10?

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

I need help solving this equation

OpenStudy (anonymous):

could you help me?

OpenStudy (anonymous):

= ($704/[1 - 1/5])/(1 - 1/12) = ($704/[5/5 - 1/5])/(12/12 - 1/12) = ($704/[4/5])/(11/12) = ($704[5/4])/(11/12) = $880(12/11) = $960 Answer: $960

OpenStudy (anonymous):

wait a minute...

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

that simplied the problem alot

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

see ya for now

OpenStudy (anonymous):

\(\Large [M - (\color {red}{M/2})] - \color {blue}{(1/5)[M - (M/2)]} = 704 \) \(\Large [M - \frac{M}{2}] - \frac{1}{5}[M - \frac{M}{2}] = 704 \)

OpenStudy (anonymous):

oops!!! 1/12... i read it as 1/2.... sorry....